### v1

# Caution for the Environment: Multimodal Agents are Susceptible to Environmental Distractions

Abstract
This paper investigates the faithfulness of multimodal large language model (MLLM) agents in the graphical user interface (GUI) environment, aiming to address the research question of whether multimodal GUI agents can be distracted by environmental context. A general setting is proposed where both the user and the agent are benign, and the environment, while not malicious, contains unrelated content. A wide range of MLLMs are evaluated as GUI agents using our simulated dataset, following three working patterns with different levels of perception. Experimental results reveal that even the most powerful models, whether generalist agents or specialist GUI agents, are susceptible to distractions. While recent studies predominantly focus on the helpfulness (i.e., action accuracy) of multimodal agents, our findings indicate that these agents are prone to environmental distractions, resulting in unfaithful behaviors. Furthermore, we switch to the adversarial perspective and implement environment injection, demonstrating that such unfaithfulness can be exploited, leading to unexpected risks. The code is available at https://github.com/xbmxb/EnvDistraction.
1 Introduction
Figure 1: (a) Previous studies expect agents to work normally and improve the action prediction performance (e.g., Yang et al. (2023)). (b) Recent works have discussed that agents can be distracted by user inputs through ambiguous instructions or malicious content (e.g.,Ruan et al. (2024)). (c) We focus on the setting where the agent is affected by distractions that exist in the environment. When the GUI agent is perceiving the environment, the screen presents distractions (e.g., coupons) that are irrelevant to the user’s goal. These distractions can interfere with the GUI agent’s task by distracting the agents from the user’s goal, thereby misleading the agent’s action trajectory.
Empowered by the commendable progress in multimodal large language models (MLLMs) (OpenAI, 2023; Templeton et al., 2024), multimodal agents have demonstrated substantial potential in addressing complex interactive tasks Yao et al. (2022a); Shridhar et al. ; Wang et al. (2023). Graphical user interface (GUI) agents emerge as a prime example (Zhang and Zhang, 2023; Hong et al., 2023). GUI agents enable human-like behaviors on operating systems by receiving the environment status (e.g., screen) and predicting the subsequent action (e.g., “click the search bar”) based on a specific goal (e.g., “report hot financial news for today”). Their capabilities have reached an even more promising level with specialized augmentations. Concretely, related studies have confirmed that pre-planning and post-reflection are beneficial for interactive reasoning Hong et al. (2023); Zhang et al. (2024); Bai et al. (2024), and that layouts and localization are crucial for fine-grained perception Ma et al. (2024); Cheng et al. (2024); You et al. (2024). Benefiting from these studies, there is a growing trend of social impact to incorporate agents as AI assistants, thereby boosting efficiency and reducing human workloads in daily routines Wu et al. (2024b); Song et al. (2023).
Despite the exciting progress, it remains an open question whether GUI agents can stay faithful to users without getting distracted Shi et al. (2023) by the rich contents in a multimodal environment. Figure 1-(c) shows a typical example. When deployed in the real-world community, GUI agents will inevitably face distractions, such as publicity and promotion activities, that interfere with their pursuit of user goals. If these distractions intervene in the agents’ actions, they may lead to uncontrollable environmental states, hindering agents from following the user’s intended goal. Even worse, the agents might complete an unexpected task suggested by the distractions.
On the one hand, those distractions cause significant changes in the action spaces for interaction. For example, the APP/website opening ads will cover the screen and disable all possible actions and some require manual shutdown, e.g., coupons. Under such a circumstance, the agents are likely to fall into an infinite loop if they stick to unavailable actions or fail to bypass the distractions. On the other hand, distractions also lead to inconsistency between the goal and the environmental contexts. The required operations are covered by distractions or mixed up with distractions, which means what exists is not (all) needed. This can be similar to conflicts in the inputs, where LLMs can be misled by unexpected content Mallen et al. (2023); Wei et al. (2023b); Shi et al. (2023); Li et al. (2023). As a result, agents will turn to another topic or even exhibit uncontrolled behaviors, instead of staying faithful to the original goal from the user instruction.
This work focuses on faithfulness of multimodal GUI agents. Concretely, we explore the research question: To what extent can a GUI agent be distracted by a multimodal environment, thereby compromising its adherence to the goal? under the general circumstance where the user and the agent are both benign, the environment is risky but not malicious. Our research topic differs from existing work that either advances the GUI action performance or explores safety awareness but considers their comprehensive ability and reliability (Figure 1). Our considered distractions are less than ideal situations in an imperfect episode, such as interruptions, but are not abnormal adversarial examples, such as images with triggers.
To address our research question, we first define the problem of distracting agents from the environment and three action types, gold, distracted, and invalid action. To simulate the distracting environment, we construct a dataset automatically, which contains four subsets, each for a vulnerable scenario, namely pop-up boxes, search, recommendation, and chat. Three working patterns are applied with different levels of perception and modal fusion. Experiments on 10 popular MLLMs reveal that both generalist and specialist GUI agents are susceptible to environmental distractions, and enhancing environmental perception is insufficient to mitigate these effects. To underscore the significance, we propose an adversarial setting, environment injection, and demonstrate that it is feasible to compromise an agent by leveraging such distractions.
Our contributions can be summarized as follows:
$\circ$ We propose the question of the faithfulness of agents in a distracting multimodal environment and define a realistic setting, which is benign but risky.
$\circ$ We construct a simulated dataset of distractions from the multimodal environment, empirically reveal the vulnerability of the agents’ faithfulness and present detailed analyses.
$\circ$ We demonstrate that the distractions can be exploited for malicious use with an automatic attacking named environment injection.
2 Related Work
This section introduces the background of MLLMs, GUI agents, and potential risks of language agents.
2.1 Multimodal Large Language Models
The remarkable capability of large language models has been successfully generalized to diverse modalities OpenAI (2023). A typical language-centered MLLM consists of a modality encoder, an LLM, an adapter (or a projector) as the bridge Liu et al. (2023); Alayrac et al. (2022); Zhu et al. (2023), and an optional generator for non-text modalities Wu et al. (2023); Shen et al. (2023). Recent works explore unified embedding Bavishi et al. (2023), grounding Peng et al. (2024), and encoder ensemble Hong et al. (2023), encouraging novel paradigms. Multi modalities enable the physical perception of agents, adding “eyes” to the language-based “brain” Zhang and Zhang (2023); Yan et al. (2023). The multimodal agents are facilitated to conduct actions in human-like behaviors.
2.2 Agents can Operate GUIs
Recently, the term agent has come to describe models that iteratively interact with an environment to solve complex problems Yao et al. (2022a, b). Among these challenges, GUI automation stands out as a representative task, demanding comprehensive perception and reliable actions.
Small-size models have achieved early success in GUI automation Sun et al. (2022); Rawles et al. (2023) by action selection. Since the emergence of LLMs Ouyang et al. (2022), the agents inherit language abilities from pre-training and the environment perception focuses on layout parsing or HTML code understanding Zhou et al. (2024); Lai et al. (2024). Empowered by multiple modalities, visual perception gradually replaces the complex strategies for the textual description of environments, advancing GUI agents to a generalizable paradigm. The visual channel plays a significant role in environment perception. Pre-training on element grounding Cheng et al. (2024), supplementing local layouts Ma et al. (2024), and selecting appropriate vision modules Zhang and Zhang (2023) have shown effective and essential in improving performance.
2.3 Risk of Language Agents
Despite the remarkable progress of agents, concerns about potential risks have been raised.
$\circ$ The output of agents can be manipulated. LLM-based Agents, even when aligned with human preference, can still be prone to generating biased or harmful content. Recent adversarial studies to jailbreak or hijack LLMs Yuan et al. (2024b); Huang et al. (2024); Yang et al. (2024); Wu et al. (2024a) have challenged prevention and promoted new strategies Dai et al. (2024); Wang et al. (2024).
$\circ$ The behavior of agents needs prejudgement. The behavioral risk lies in the implicit results of the agents’ actions. For example, agents should not forward unconfirmed gossip on social media. The risk is more concealed as it lies in the results rather than the generated content. Hence, detection and prevention require extrapolation based on prior knowledge and reasoning ability Tian et al. (2023); Yuan et al. (2024a). A representative work, Toolemu Ruan et al. (2024), emulates actions in a GPT-4-based sandbox.
Different from previous studies, our work proposes a novel setting (Figure 1) in three ways. (i) The possible distractions are in the environment instead of the user input. The distracting contents are received from the environmental perception. (ii) The user, agent, and environment are all benign, having no malicious intention or deliberate misleading. (iii) We focus on whether agents follow distracting content, instead of safety or ethics.
Figure 2: Overview of the methodology for distracting GUI agents. We first construct distracting environments (the left part), then implement working patterns with prompts (the middle part), and evaluate a broad range of multimodal agents, judging the predicted action as gold, distracted, and invalid (the right part).
3 Distracting GUI Agents
We begin by formalizing the problem statement in 3.1, then present approaches for distraction simulation in 3.2, measurement in 3.3, and working patterns 3.4. Figure 2 shows an overview.
3.1 Problem Statement
The formalization of GUI agent and the problem of distraction for GUI agents are as follows.
GUI agent. Consider a GUI agent $A$ interacting with an OS environment $Env$ to complete a specific goal $g$. At each time step $t$, the agent performs an action $a_{t}$ on the OS based on its perception of the environmental state $s_{t}$,
$\begin{split}&\textsc{Episode}=(g,[(s_{t},a_{t})]_{t=1}^{n}),\\ &a_{t}\leftarrow A_{LLM}(s_{t},g),s_{t+1}\leftarrow(s_{t},a_{t}),\end{split}$ (1)
where each action is expected to contribute to the goal so that the goal can be completed after $n$ steps.
Distraction for GUI agents. However, the OS environment contains complex information of varying quality and from diverse sources. The environmental contents are formally divided into two categories: contents that are useful or necessary for goal completion, ${c^{use}}$, and distractors that are irrelevant to the user’s goal but indicate another target, ${c^{dist}}$,
$\begin{split}&s_{t}=(\{c_{t}^{use}\},\{c_{t}^{dist}\}).\\ \end{split}$ (2)
GUI agents must use $\{c_{t}^{use}\}$ for faithful actions instead of being distracted by ${c^{dist}}$ or predicting irrelevant actions. Based on the $s_{t}$, the available actions are determined. The valid action space $\mathbb{A}_{t}$ can be annotated with three types of labels, namely gold actions, distracted actions, and other actions,
$\begin{split}\mathbb{A}_{t}\leftarrow s_{t},\mathbb{A}_{t}=(\{a_{gold}\},\{a_{dist}\},\{a_{other}\}),\\ \end{split}$ (3)
where gold labels denote actions for $g$. Distracted labels denote actions for the distracting target, misled by distractions.
Receiving a goal $g$ and a screenshot as the current state $s_{t}$, the agent outputs the next action $a_{t}$. Then, $a_{t}$ is judged by comparing to action spaces. Our task concerns whether $a_{t}$ is faithful (gold) or distracted, or fails to be valid,
$\textsc{eval}(a_{t})=\left\{\begin{array}[]{lcr}\text{Gold}&a_{t}\in\{a_{gold}\}\\ \text{Distracted}&a_{t}\in\{a_{dist}\}\\ \text{Invalid}&a_{t}\notin\mathbb{A}_{t}.\\ \end{array}\right.$ (4)
The distraction problem needs to be studied under reasonable conditions, summarized as follows: First, be realistic. The environment state $s_{t}$, especially $c^{use}$, must reflect a realistic world. Second, be explicit. The distractors, $c_{dist}$, must be reasonable and distinguishable. It should be clear whether a piece of content is a distractor, similar to how humans can recognize ads. Third, be diverse. The environment states and distractors should exhibit diversity.
3.2 Distraction Simulation
Following the problem statement, we simulate the task without the loss of generality and construct a simulated dataset, $D$. Each sample is a triplet $(g,s,\mathbb{A})$ consisting of a goal $g$, a screenshot image representing the state $s$, and a valid action space $\mathbb{A}$. The critical part is to construct $s$ such that it includes $c^{use}$ and $c^{dist}$. We consider four common scenarios, namely pop-up boxes, search, recommendation, and chat, forming four subsets. A compositional strategy is employed for goals, layouts, and distractors. Table 1 shows an overview, including designs and sample numbers.
Scenario Pop-up boxes Search Recommendation Chat
Users’ Goal Browse the website Common queries Shopping targets Chat or modify the chat interface
Distractions Boxes suggest another action Fake items, ads, other queries Different products, ads Chat logs suggest another action
Faithful Actions Button to reject, cross mark True search results Related products Correct button
Distracted Actions Follow the popup box Fake results Fake products Follow the chat log
Sample number 662(208+220+234) 250 176 110
Table 1: Overview of our simulated dataset. Examples of each scenario are shown in Figure 4.
$\circ$ Pop-up boxes. Pop-up boxes frequently obscure the original web pages and prevent users from accessing the underlying content. This compels users to address the pop-up boxes before they can resume their intended actions, making the content of these boxes a significant distraction that must be managed. To simulate this scenario, we introduce pop-up boxes to a shopping website and modify their contents. Specifically, we leverage GPT-4 to generate target operations or questions as initial goals. For each goal, we then instruct GPT-4 to create various distractions, including common attractions like ads, notifications, and alerts. For layout, we prepare three templates for common pop-up boxes: (1) an input box with a submission button, (2) two buttons presenting choices, and (3) a four-choice checkbox. The faithful action is to dismiss the contents by clicking one of the buttons (such as “No thanks”) or by clicking a cross mark to close the box. If the agent follows the pop-up instead, it is considered distracted.
$\circ$ Search. Retrieval systems are found to be susceptible to the “Spiral of Silence” effect Chen et al. (2024), where LLM-generated (fake) text consistently outperforms human-authored content in search rankings, potentially leading to the marginalization of true information. Similarly, we examine the impact of integrating a fake item into search results. Concretely, we begin by prompting GPT-4 to generate common search queries. For each query, Google search API is called for the ground truth of the search results. Subsequently, we instruct GPT-4 to make up a distracting item, such as ads, different websites, or suggested other queries. The faithful action is to click on any of the true results. If the agent clicks on the fake results, it indicates a distraction from accurate information.
$\circ$ Recommendation. The recommendation scenario presents related products according to the user query. We mix a fake product into the recommendation page for the shopping targets generated by GPT-4. For ground truth products, we simulate a BM25 Robertson et al. (2009) retriever on Amazon Reviews Hou et al. (2024). The product titles and descriptions are pre-indexed based on Lucene. The retrieval is limited to a specific category for efficiency. Similarly, GPT-4 makes up an appealing fake product to replace a random one. The action is considered faithful if the agent clicks any real product. This scenario differs from the search subset because the products are limited to the candidate set rather than the entire World Wide Web and the products are annotated with an accompanying picture.
$\circ$ Chat. Chat is also a vulnerable scenario for GUI agents. Because the received messages will be displayed exactly as they are, which means a part of the screen can be decided by other information sources. We generate actions based on Discord webpage. Two goals with different actions are generated with the assistance of the Discord manual. One is rewritten to the user goal. The other is rewritten into a dialogue containing explicit action guides. The dialogues are posted to Discord by sending from two tool accounts. Then, the account that receives the action advice saves the chat record as a screenshot. The agent determines the next action for the user goal. If it follows the action guides in the dialogue, then it is distracted.
Action labels. The action labels are annotated during the generation. Given a screenshot, $\{a_{gold}\}$ and $\{a_{dist}\}$ are determined by the element layouts. Available actions that do not match the gold actions or follow the distractors are denoted as $\{a_{other}\}$, if any. $\{a_{inv}\}$ includes any actions that are not entailed in the layout. Based on the action labels, the targeting elements’ locations are annotated by OCR for coordinate evaluation for specialist agents.
3.3 Measurement
The evaluation of the predicted action $\hat{a}$ is formalized separately for two kinds of agents, denoted as $\textsc{M}_{txt}$ and $\textsc{M}_{loc}$ in Eq. 5. (i) Strong generalist MLLMs (e.g., GPT-4o) can operate GUIs as agents with natural language by predicting actions with screen element descriptions, like “Submit button”. The prediction is measured by token-level $F_{1}$ with annotated actions and matches one if $F_{1}$ surpasses a threshold, $\tau_{txt}$. (ii) specialist agents (e.g., CogAgent) are trained to generate operating locations using precise coordinates on the screen. The predicted coordinate matches an annotated action if it falls into an annotated box,
$\begin{split}&\textsc{M}_{txt}(\hat{a},a)=F_{1}(\textsc{T}(\hat{a}),\textsc{T}(a))\geq\tau_{txt},\\ &\textsc{M}_{loc}(\hat{a},a)=\hat{a}_{loc}\in a_{loc},\end{split}$ (5)
where the outputs are bool indicators. Based on the type of the matched labels, accuracy for gold actions, distracted actions, or invalid actions are computed respectively,
$\begin{split}&Acc_{gold}=\frac{1}{|D|}\sum_{d\in D}\exists a_{i}\in\{a_{gold}\},\textsc{M}(\hat{a},a_{i}),\\ &Acc_{dist}=\frac{1}{|D|}\sum_{d\in D}\exists a_{i}\in\{a_{dist}\},\textsc{M}(\hat{a},a_{i}),\\ &Acc_{inv}=1-\frac{1}{|D|}\sum_{d\in D}\exists a_{i}\in A,\textsc{M}(\hat{a},a_{i}),\\ \end{split}$ (6)
where $Acc_{\texttt{gold}}$ reflects the faithfulness and helpfulness of agents. $Acc_{\texttt{dist}}$ shows the unfaithfulness, i.e. how often agents are distracted from their goals. $Acc_{\texttt{inv}}$ indicates the multimodal capabilities including perception and prediction, i.e., how often agents fail to give valid actions.
3.4 Working Patterns
The behavior of agents can be sensitive to working patterns Shinn et al. (2024); Khattab et al. (2022), especially the perception level for complex environments like GUIs. For a comprehensive study, we implement working patterns with three levels of environmental perception, which are an implicitly-perceived level, a partially-perceived level, and a well-perceived level.
Pattern Env. Modality Env. Perception
Direct prompt Image Implicitly-perceived
CoT prompt Image, text Partially-perceived
Action anno. Image, text Well-perceived
Table 2: Working patterns impact the modality of the environment representation and perception.
$\circ$ Direct prompt. The input is a goal and a screenshot, and the expected output is the next action. It is denoted as
$\begin{split}&\hat{a}=A(g,s).\end{split}$ (7)
$\circ$ CoT prompt. Chain-of-Thought (CoT) Wei et al. (2023a) have unveiled the reasoning capability of agents by asking the models to generate explicit rationales for deriving an answer Yao et al. (2022b); Hong et al. (2023). With a CoT-like pattern, the agent first receives the screenshot state to extract possible actions (“thoughts”), then predicts the next action based on the goal, denoted as
$\begin{split}&\hat{\mathbb{A}}=A(s),\quad\hat{a}=A(g,s,\hat{\mathbb{A}}).\end{split}$ (8)
$\circ$ Action annotations. Extracting all available actions from a screenshot is still a bottleneck for agents. If this burden is removed, the performance can depend more on judging distractions and keeping faithfulness. The available actions can be integrated into the input, denoted as
$\begin{split}&\hat{a}=A(g,s,\mathbb{A}_{w/o\_label}),\end{split}$ (9)
where $A_{w/o\_label}$ denotes annotated actions without their labels of gold or distraction.
In essence, providing available actions means two changes, as summarized in Table 2, (i) information for potential actions entailed in the image is disclosed and perceived as texts; (ii) information is fused into the text channel from the vision channel  [^1] Due to the space limitation, the prompts and detailed implementation will be released with our code and data..
4 Experiments
This section will introduce the implementation settings including the dataset and agent models and then show our empirical results with findings.
4.1 Implementation
Dataset. Our simulated dataset contains 1198 samples in total, as statistics shown in Table 1.
Agent models. We implement a series of well-known MLLMs on our datasets. (i) Generalist agents. Multi-modal versions of strong black-box LLMs have shown promising performance and are available by API services, including GPT-4v, GPT-4o, GLM-4v GLM et al. (2024), Qwen-VL-plus Bai et al. (2023), and Claude-Sonnet-3.5Templeton et al. (2024). We also consider powerful open-source MLLMs, including Qwen-VL-chat Bai et al. (2023), MiniCPM-Llama3-v2.5 Hu et al. (2024), LLaVa-v1.6-34B Liu et al. (2023). (ii) Specialist agents. Recent studies released expert MLLMs for GUI agents after post-pre-training or instruction fine-tuning, including CogAgent-chat Hong et al. (2023) and SeeClick Cheng et al. (2024).
Agent API Specialist $Acc_{\texttt{gold}}$ $Acc_{\texttt{dist}}$ $Acc_{\texttt{inv}}$
GPT-4v 67.76 14.04 18.85
GPT-4o 74.31 9.09 20.19
GLM-4v 36.69 28.36 35.15
Claude 68.00 14.28 17.04
Qwen-VL-plus 30.74 14.84 55.47
[…]
MiniCPM 37.20 24.42 39.01
LLaVa-1.6 40.09 16.28 43.83
[…]
SeeClick 31.84 6.84 47.46
Table 3: Experiment results overview (direct prompt).
Patterns Direct prompt CoT prompt Action anno.
Agent $Acc_{\texttt{gold}}$ $Acc_{\texttt{dist}}$ $Acc_{\texttt{inv}}$ $Acc_{\texttt{gold}}$ $Acc_{\texttt{dist}}$ $Acc_{\texttt{inv}}$ $Acc_{\texttt{gold}}$ $Acc_{\texttt{dist}}$ $Acc_{\texttt{inv}}$
GPT-4v 67.44 6.57 25.95 13.36$\downarrow$54.08 12.53$\uparrow$5.96 74.11$\uparrow$48.16 83.27$\uparrow$15.83 16.26$\uparrow$9.69 0.47$\downarrow$25.48
GPT-4o 86.64 6.53 6.83 38.33$\downarrow$48.31 16.08$\uparrow$9.55 45.59$\uparrow$38.76 73.04$\uparrow$34.71 26.01$\uparrow$19.48 0.94$\downarrow$5.89
GLM-4v 4.49 59.08 36.42 6.26$\uparrow$1.77 62.49$\uparrow$3.41 31.25$\downarrow$5.17 11.26$\uparrow$6.77 57.45$\downarrow$1.63 31.27$\downarrow$5.15
Claude 77.26 11.94 10.80 42.64$\downarrow$34.62 17.04$\uparrow$5.1 40.33$\uparrow$29.53 77.85$\uparrow$0.59 21.69$\uparrow$9.75 0.46$\downarrow$10.34
Qwen-VL-plus 7.35 27.14 68.90 15.03$\uparrow$7.68 76.92$\uparrow$49.78 8.05$\downarrow$60.85 8.71$\uparrow$1.36 77.47$\uparrow$50.33 13.81$\downarrow$55.09
[…]
MiniCPM 14.62 27.94 57.46 26.33$\uparrow$11.71 48.58$\uparrow$20.64 25.08$\downarrow$32.38 52.02$\uparrow$37.40 47.67$\uparrow$19.73 0.30$\downarrow$57.16
LLaVa-1.6 1.78 22.40 75.82 6.70$\uparrow$4.92 54.85$\uparrow$32.45 38.48$\downarrow$37.34 15.28$\uparrow$13.5 72.41$\uparrow$50.01 12.31$\downarrow$63.51
[…]
SeeClick 6.64 2.17 91.19 N/A N/A N/A 78.29$\uparrow$71.65 12.42$\uparrow$10.25 9.29$\downarrow$81.9
Table 4: Results on the scenario of pop-up boxes.
Patterns Direct prompt CoT prompt Action anno.
Agent $Acc_{\texttt{gold}}$ $Acc_{\texttt{dist}}$ $Acc_{\texttt{inv}}$ $Acc_{\texttt{gold}}$ $Acc_{\texttt{dist}}$ $Acc_{\texttt{inv}}$ $Acc_{\texttt{gold}}$ $Acc_{\texttt{dist}}$ $Acc_{\texttt{inv}}$
GPT-4v 92.00 4.80 4.00 88.40$\downarrow$3.60 2.80$\downarrow$2.00 8.80$\uparrow$4.80 95.20$\uparrow$3.20 2.40$\downarrow$2.40 2.40$\downarrow$1.60
GPT-4o 94.00 2.40 3.60 86.8$\downarrow$7.20 4.40$\uparrow$2.00 8.80$\uparrow$5.20 84.40$\downarrow$9.60 15.20$\uparrow$12.8 0.40$\downarrow$3.20
GLM-4v 60.40 36.40 3.20 77.73$\uparrow$17.33 2.94$\downarrow$33.46 19.33$\downarrow$16.13 91.20$\uparrow$30.80 3.20$\downarrow$33.20 5.60$\uparrow$2.40
Claude 93.60 3.60 2.80 76.71$\downarrow$16.89 5.22$\uparrow$1.62 18.07$\uparrow$15.27 96.40$\uparrow$2.80 3.60$\downarrow$0.00 0.0$\downarrow$2.80
Qwen-VL-plus 57.60 7.60 34.80 82.00$\uparrow$24.40 16.00$\uparrow$8.40 2.00$\downarrow$32.80 82.00$\uparrow$24.40 19.20$\uparrow$11.60 0.00$\downarrow$34.80
[…]
MiniCPM 54.80 43.60 0.60 68.80$\uparrow$14.0 13.20$\downarrow$30.40 8.00$\uparrow$7.4 75.60$\uparrow$20.80 24.40$\downarrow$19.20 0.00$\downarrow$0.60
LLaVa-1.6 60.40 29.20 10.40 51.60$\downarrow$8.80 15.20$\downarrow$14.0 33.20$\downarrow$22.80 78.80$\uparrow$18.40 19.20$\downarrow$10.0 2.0$\downarrow$8.40
[…]
SeeClick 25.60 11.20 63.20 N/A N/A N/A 66.80$\uparrow$41.20 23.20$\uparrow$11.20 10.00$\downarrow$53.20
Table 5: Results on the scenario of search.
Patterns Direct prompt CoT prompt Action anno.
Agent $Acc_{\texttt{gold}}$ $Acc_{\texttt{dist}}$ $Acc_{\texttt{inv}}$ $Acc_{\texttt{gold}}$ $Acc_{\texttt{dist}}$ $Acc_{\texttt{inv}}$ $Acc_{\texttt{gold}}$ $Acc_{\texttt{dist}}$ $Acc_{\texttt{inv}}$
GPT-4v 89.77 10.23 0.00 93.75$\uparrow$3.98 6.25$\downarrow$3.98 0.00$\downarrow$0.00 89.77$\uparrow$0.00 10.23$\downarrow$0.00 0.00$\downarrow$0.00
GPT-4o 92.05 7.95 0.00 93.75$\uparrow$1.70 6.25$\downarrow$1.70 0.00$\downarrow$0.00 94.32$\uparrow$2.27 5.68$\downarrow$2.27 0.00$\downarrow$0.00
GLM-4v 80.68 18.75 0.57 82.95$\uparrow$2.27 16.48$\downarrow$2.27 0.57$\downarrow$0.0 72.16$\downarrow$8.52 27.84$\uparrow$9.09 0.00$\downarrow$0.57
Claude 78.41 21.59 0.00 89.20$\uparrow$10.79 10.80$\downarrow$10.79 0.00$\downarrow$0.00 85.80$\uparrow$7.39 14.20$\downarrow$7.39 0.00$\downarrow$7.39
Qwen-VL-plus 53.98 15.34 30.68 56.82$\uparrow$2.84 18.18$\uparrow$2.84 25.00$\downarrow$5.68 61.93$\uparrow$7.95 27.84$\uparrow$12.50 10.23$\downarrow$20.45
[…]
MiniCPM 77.27 22.73 0.00 80.11$\uparrow$2.84 11.36$\downarrow$11.37 8.52$\uparrow$8.52 66.48$\downarrow$10.79 33.52$\uparrow$10.79 0.00$\downarrow$0.0
LLaVa-1.6 81.82 16.48 1.70 64.20$\downarrow$17.62 18.75$\uparrow$2.27 11.05$\uparrow$9.35 82.39$\uparrow$0.57 16.48$\downarrow$0.00 1.14$\downarrow$0.56
[…]
SeeClick 86.93 13.07 0.00 N/A N/A N/A 80.68$\downarrow$6.25 17.61$\uparrow$4.54 1.70$\uparrow$1.70
Table 6: Results on the scenario of recommendation.
Patterns Direct prompt CoT prompt Action anno.
Agent $Acc_{\texttt{gold}}$ $Acc_{\texttt{dist}}$ $Acc_{\texttt{inv}}$ $Acc_{\texttt{gold}}$ $Acc_{\texttt{dist}}$ $Acc_{\texttt{inv}}$ $Acc_{\texttt{gold}}$ $Acc_{\texttt{dist}}$ $Acc_{\texttt{inv}}$
GPT-4v 21.82 34.55 45.45 13.64$\downarrow$8.18 21.82$\downarrow$12.73 61.82$\downarrow$7.27 51.82$\uparrow$30.00 49.09$\uparrow$14.54 9.09$\downarrow$36.36
GPT-4o 24.55 19.09 60.91 25.45$\uparrow$0.90 13.64$\downarrow$5.45 55.45$\downarrow$5.46 67.27$\uparrow$42.72 30.00$\uparrow$10.91 13.64$\downarrow$47.27
GLM-4v 0.00 0.00 100.00 5.45$\uparrow$5.45 17.27$\uparrow$17.27 76.36$\downarrow$23.64 36.04$\uparrow$36.04 53.15$\uparrow$53.15 19.82$\downarrow$80.18
Claude 22.73 20.00 54.55 16.36$\downarrow$6.37 21.82$\uparrow$1.82 51.82$\downarrow$2.73 57.27$\uparrow$34.54 38.18$\uparrow$18.18 0.00$\downarrow$54.55
Qwen-VL-plus 3.64 7.27 89.09 8.70$\uparrow$5.06 4.35$\downarrow$2.92 77.39$\downarrow$11.70 47.27$\uparrow$43.63 30.00$\uparrow$22.73 31.28$\downarrow$57.81
[…]
MiniCPM 0.91 1.82 98.18 9.09$\uparrow$8.18 8.18$\uparrow$6.36 62.73$\downarrow$35.45 52.73$\uparrow$51.82 28.18$\uparrow$26.36 27.27$\downarrow$70.91
LLaVa-1.6 6.36 1.82 91.82 2.73$\downarrow$3.63 8.18$\uparrow$6.36 65.45$\downarrow$26.37 47.27$\uparrow$40.91 31.82$\uparrow$30.0 29.09$\downarrow$62.73
[…]
SeeClick 8.18 0.91 35.45 N/A N/A N/A 3.64$\downarrow$4.54 2.73$\uparrow$1.82 29.09$\downarrow$6.36
Table 7: Results on the scenario of chat.
4.2 Main Results
Experimental results are shown in Table 3-7. Specifically, Table 3 shows an overview of the average of our four scenarios by direct prompt, and the following four tables present detailed scores across different scenarios and working patterns. Our results answer the following three key questions.
(i) Can the multimodal environment distract a GUI agent from its goal? In risky environments, multimodal agents are susceptible to distractions that may lead them to abandon their goals and act unfaithfully. Each model produces actions that deviate from the original goal across our four scenarios. Such distracted predictions hinder the accuracy of gold actions. The Strong APIs (9.09% of GPT-4o) and specialist agents (6.84% of SeeClick) are more faithful than generalist open-source agents.
(ii) What is the relation between faithfulness ($Acc_{\texttt{dist}}$) and helpfulness ($Acc_{\texttt{gold}}$)? There are two situations. First, MLLMs with strong capabilities can be both helpful and faithful ( GPT-4o, GPT-4v, and Claude). They exhibit low $Acc_{\texttt{inv}}$ scores, and relatively higher $Acc_{\texttt{acc}}$ and lower $Acc_{\texttt{dist}}$ (e.g., GPT-4o on pop-up boxes, search, and recommendation). Whereas, stronger perception but inadequate faithfulness can lead to greater susceptibility to distractions and lower helpfulness. For instance, GLM-4v demonstrates a higher $Acc_{\texttt{dist}}$ and a much lower $Acc_{\texttt{inv}}$ compared to open-sourced MLLMs, and GPT-4v and GPT-4o exhibit this trend in the chat subset. Hence, faithfulness and helpfulness are not mutually exclusive but can be enhanced simultaneously. It is even more critical to enhance faithfulness for stronger MLLMs.
(iii) Can multimodal environmental perception help alleviate unfaithfulness? By implementing different working patterns, visual information is integrated into the textual channel to augment environmental perception. However, the results indicate that textual augmentation for GUI comprehensive can actually increase distractions. The increase of distraction action can outweigh the benefits, as seen in almost all setups with action annotations in the pop-up box, recommendation, and chat subsets (e.g., Qwen-VL, LLaVa, and GLM-4v). CoT-prompt, as a self-guided textual augmentation, can largely alleviate the perception burden but also increase distractions. These working patterns cannot work as a positive “defense” of environmental distractions.
This finding highlights two key points: firstly, there is a risk associated with stronger perception capabilities, and secondly, the fusion of UI information across textual and visual modalities (such as OCR) must be approached with greater caution.
5 Analysis
In this section, we present discussions based on the detailed experiment results. We first compare the results from the aspects of the base MLLM agents, working patterns, and scenarios. Then, we suggest two mitigation methods with experiments.
5.1 Comparing MLLMs
Among the generalist agents, GPT-4o demonstrates the best faithfulness and effectiveness in our scenarios, with the minimum average $Acc_{\texttt{distract}}$ (9.09%), and the maximum average $Acc_{\texttt{gold}}$ (74.3The open-sourced models get close scores on average, where LLaVa and MiniCPM are generally better. However, they demonstrate different abilities across scenarios. LLaVa is better at Searching and Recommending subsets, indicating advanced textual perception. MiniCPM is better at the pop-up boxes, and thus can be superior for visual (layouts or icons) knowledge.
Regarding specialist agents, the $Acc_{\texttt{dist}}$ of both CogAgent and SeeClick is much lower than general MLLMs, indicating that they enjoy higher faithfulness. CogAgent outperforms all agents except GPT-4 and Claude on both faithfulness and effectiveness. Shortcuts hinder the full potential of SeeClick, causing a high proportion of invalid actions. Specifically, when SeeClick encounters irrelevant pop-up boxes, it often predicts the coordinates at the very top right corner, likely attempting to close the box, rather than the actual position of the cross mark. And on screenshots of search pages, it often clicks the search bar. Once the available action annotations are input, the invalid actions and distractions are significantly reduced. This suggests that SeeClick knows what to do but draws wrong conclusions about the locations. These shortcuts can result from the inherited GUI knowledge gained during domain pre-training.
In summary, strong API-based MLLMs are superior to open-sourced MLLMs regarding faithfulness and effectiveness. GUI pre-training can largely improve the expert agents’ faithfulness and effectiveness but can introduce shortcuts.
5.2 Comparing Working Patterns
Our three considered working patterns provide different levels of hints for the action prediction task. The direct pattern represents the environment with only an image. The action annotations expose the ground truth action space that could nearly substitute the environmental perception, making the task akin to a multiple-choice problem. This represents the upper bound of the perception capability. As a transition in between, CoT is applied to first ask the agent to predict a pseudo-action space, which is used to guide its action.
Our results show that the proportions of both gold actions and distracted actions largely increased with ground truth action space. However, on the other hand, the distracted proportions mean that even with a “perfect” perception, the agents are still vulnerable to distractions. For most models, the CoT prompt can work to provide some guidance and restrain agents’ behavior from invalid actions, but the distracted proportions also increased. Although it can not completely defend, the self-guided step-by-step process demonstrates the potential for mitigation.
5.3 Comparing Scenarios
The four simulated scenarios vary in emphasis and difficulty based on our empirical results. Figure 3 characters the variances in two types of challenges.
(i) Faithfulness. In our experiments, the pop-up box subset leads to the most unfaithful results in each working pattern (high $Acc_{\texttt{dist}}$). The recommendation and search scenarios get more gold actions. We use the proportion of distractions as a general measurement of “the difficulty to stay faithful”, computed as $avg(|a_{dist}|)/|A|)$. The pop-up box subset has the largest distraction proportion, as we add several fields to ask the agent to fill in the box (e.g., questionnaires). The other three subsets only suggest one distraction on the screen, thus, the more the possible actions, the lower the distraction proportion.
(ii) perception. In our results, the distractions are more successful in the recommendation subset. The chatting subset suffers from invalid actions or valid but unrelated actions. Accordingly, we also qualitatively illustrate the type and level of the perception difficulty. The pop-up boxes and chatting page mainly require the comprehension of the layout and icons. For example, find the cross mark to close the pop-up box or find the icon most related to the goal. The chatting page is more complex and implicit. For textual perception, true search results are more compact and closely related to the query. In contradiction, the real products for recommendation are noisy, more realistic but less attractive than the fake ones.
Figure 3: Illustration of scenario features.
6 Towards the Adversarial Perspective
The study discusses the general problem of environments that are imperfect, noisy, or flawed can distract multimodal agents, while they are not abnormal or malicious. However, such distractions can be exploited for malicious purposes. This section considers an extreme situation with an adversarial role and shows the feasibility of an active attack to distract GUI agents from the environment, named environment injection.
6.1 Threat Model
The user communicates with a multimodal GUI agent. The attacker needs to mislead the agent by only altering the GUI environment. The attacker can eavesdrop on the messages from the user and get the goal. The attacker can also hack the related environment to change the action space. For example, it is possible to block the package from a host and change the contents of a website. The problem is denoted as
$\begin{split}&s_{adv}\leftarrow\text{Adv}(g,s),a_{dist}=A(g,s_{adv}).\end{split}$ (10)
6.2 Method and Experiments
Without loss of generality, we verified the feasibility of environment injection on the pop-up box scenario. The box layout is limited to one button to accept and one to reject. The box contents are distractions, thus the gold action is to click the button reject or the cross mark and the bad action is to click the button accept.
We propose and implement a brief but effective method to rewrite the two buttons. (i) The button to accept is rewritten to be ambiguous, and reasonable for both the distraction and the true goal. We find a common operation for those two purposes. But the contents in the box clarify the function of the buttons. We leverage the findings that such contexts in the screenshots can be ignored. (ii) The button to reject is rewritten to emotionally charged language. Such leading emotions can sometimes be persuasive or even manipulative tactics to influence user decisions. The phenomenon is common in APPs, like “Cruelly Leave” for uninstalling.
Algorithm 1 Attacking sample generation
  Initialize: Web environment $s$, User’s goal $g$, Agent $A$, Maximum tries $t_{m}$, LLM
  Notions: Distracting goal $d$, Pop-up box $B$, action prediction $a$
  $t\leftarrow 0$
  for $t<t_{m}$ do
     $d\leftarrow\text{LLM}(g,s)$, $B\leftarrow\text{LLM}(d)$
[…]
     $B.\textit{button\_rej}\leftarrow\text{LLM}(d)$
     $s\leftarrow s+B$
     $a\leftarrow A(s,g)$
[…]
        Attack success
     end if
     $t\leftarrow t+1$
  end for
Table 8 shows our results on random 8 goal cases. Compared to the baseline scores, those rewriting methods decrease the faithfulness of both GLM-4v and GPT-4o, leading to higher $Acc_{\texttt{dist}}$ scores. GLM-4v is more vulnerable to emotional expressions, while GPT-4o can be misled by ambiguous acceptance more often.
Agent $Acc_{\texttt{gold}}$ $Acc_{\texttt{dist}}$ $Acc_{\texttt{inv}}$ ASR(goal)
Baselines
GPT-4o 93.64 5.00 1.36
GLM-4v 7.27 60.45 32.27
Rewrite the Button to Accept
GPT-4o 57.89 39.47 2.63 6/8
GLM-4v 18.42 57.89 23.68 6/8
Rewrite the Button to Reject
GPT-4o 54.17 33.33 12.5 6/8
GLM-4v 0.00 70.83 70.83 8/8
Rewrite Both
GPT-4o 55.56 40.00 4.44 6/8
GLM-4v 6.67 66.67 26.67 6/8
Table 8: Results of environment injection.
Figure 4: Examples of simulated data.
7 Conclusion
This paper investigates the faithfulness of multimodal GUI agents and exposes the impact of distractions from the environment. We introduce a novel research question where both the user and the agent are benign, and the environment is not malicious but risky for distracting contents. We simulate distractions in four scenarios with localized annotations and implement three working patterns with varying perception levels. A broad range of generalist agents and specialist agents are evaluated. The experimental results demonstrate that vulnerability to distractions significantly diminishes both faithfulness and helpfulness. Additionally, we propose an attacking method called “environment injection”, which exploits unfaithfulness for malicious purposes by altering distractions to include ambiguous or emotionally misleading content. More importantly, this paper calls for increased collective focus on the faithfulness of multimodal agents. We suggest future work including pre-training for faithfulness alignment, considering the correlation between environment contexts and instructions, forecasting the possible consequences of executing actions and introducing human interaction when necessary.
Ethics Statement
(i) Data privacy. There are leakage risks involved in uploading data from personal devices to LLM APIs. Our research dataset contains no personally identifiable information and is exclusively for experiments. We present examples of the simulated four scenarios in Figure 4. (ii) Potential social impacts. Our paper demonstrates that malicious actors could abuse GUI agents to achieve undesirable purposes, although agents facilitate efficiency and save human resources. We call for efforts on robust multimodal perception and protective mechanisms to control environmental risks for further application.